# Context as Memory: Scene-Consistent Interactive Long Video Generation with Memory Retrieval

### v1

@@ Abstract
Figure 1: We propose Context-as-Memory, which leverages historical frames to ensure scene consistency in long video generation by guiding the synthesis of new frames. Key frames from generated videos along camera trajectories are shown, with camera poses marked in red. It can be observed that different frames maintain good consistency when viewing the same scene from different viewpoints.
11footnotetext: Corresponding authors. Work done during an internship at KwaiVGI, Kuaishou Technology.
1 Introduction
@@
Recent breakthroughs in video generation models [28, 32, 47, 39, 22] have shown remarkable progress. Due to their powerful generative capabilities developed through training on large-scale real-world datasets, these models are considered to have the potential to become world models capable of modeling reality [28, 46, 45, 30]. Among various research directions in this field, interactive long video generation has emerged as a crucial one since many applications, such as gaming [6, 38, 52] and simulation [17, 11, 33], require interactive long video generation, where the videos are generated in a streaming manner controlled by user interactions. Recent works on long video generation [4, 34, 21, 40, 8, 12, 54] have facilitated research in this field.
@@
Despite these advances, current approaches still face significant challenges in terms of memory capabilities [51, 50], which refers to a model’s ability to maintain content consistency during continuous video generation, such as preserving the scene when the camera returns to a previously viewed location. Take Oasis [5] as an example: while it can generate lengthy Minecraft gameplay videos, even simple operations like turning left and then immediately right result in completely different scenes. This issue is prevalent across various state-of-the-art methods [38, 18, 34, 52], suggesting that while current approaches can generate videos of extended duration, they struggle to maintain coherent long-term memory of scene content and spatial relationships.
@@
In our view, these methods’ limitations in memory capabilities are not surprising. This is because when generating each new video frame, these methods can only predict based on a limited number of previous frames. For instance, Diffusion Forcing [4, 34] can only utilize context from a fixed window of several dozen frames. While this setup works for video continuation, it fails to maintain long-term consistency. In the case of video generation, if each frame to be generated could reference all previously generated frames, the generative model could select and replicate relevant content from historical frames into the current frame being generated, thus it would be possible to maintain scene consistency in long videos. In other words, all previously generated context frames serve as the memory.
@@
However, the idea of “all historical context as memory" seems intuitive but is impractical for three main reasons: (1) Including all historical frames in computation would be extremely resource-intensive. (2) Processing all historical frames is computationally wasteful since only a small fraction is relevant to the current frame generation. (3) Processing irrelevant historical frames adds noise that may hinder rather than help current frame generation. Therefore, a reasonable approach is to retrieve a small number of relevant frames from historical context as conditions for current generation, which we call “Memory Retrieval".
@@
  • We propose Context-as-Memory, highlighting the storage of frames as memory and conditioning via historical context learning for scene-consistent video generation.
  • To utilize relevant history frames efficiently, we design Memory Retrieval, a rule-based approach using FOV overlap of camera trajectory.
@@
  • Our experiments show superior long video generation memory, significantly outperforming SOTAs and achieving memory even in unseen, open-domain scenarios.
2 Related Work
2.1 Interactive Long Video Generation
@@
Video Generation Model. Video generation models can generate video sequences $\mathbf{x}=\{x^{0},x^{1},\dots,x^{t}\}$, where $x^{i}$ indicates the $i$-th frame. The current mainstream architecture is based on diffusion models [16, 36, 35, 24, 25], which excel in generating high-quality content and have been widely adopted [28, 47, 20, 7, 32, 3, 39, 22]. Other alternative architectures include next-token prediction [44, 21, 40] and various hybrid approaches [4, 8, 23].
@@
Controllable Video Generation. This task can be formulated as $p(\mathbf{x}|c)$, where $c$ represents different types of control signals. The most representative control signals include: camera motion control [41, 14, 10, 2, 1], and agent action control in games or simulators [38, 6, 52, 9, 5]. These control signals greatly enhance user interactive experience, enabling free exploration in the created virtual worlds.
@@
Streaming Video Generation. Streaming video generation can condition on previously generated frames to continuously generate new video frames, which can be expressed as $p(x^{0},x^{1},...,x^{n})=\prod_{i=0}^{n}p(x^{i}|x^{0},x^{1},...,x^{i-1})$, where $x^{i}$ indicates the $i$-th frame. Representative approaches include Diffusion-based methods [4, 34, 52, 12, 54] and GPT-like next token prediction methods [21, 40, 18]. Diffusion-based methods generally achieve higher visual quality and faster sampling speed, thus we focus on diffusion models for long video generation in this work. Although these SOTA methods generally fail to generate long videos with scene-consistent memory, instead only producing long videos with short-term continuity.
@@
Memory Capability for Video Generation. Many related works’ demos [38, 5, 34, 18] have shown that current long video generation methods generally lack memory capability: while maintaining frame-to-frame continuity, the scenes continuously change. One potential approach [49, 53, 26, 31] is to leverage 3D reconstruction to build explicit 3D representations from generated videos, then render initial frames from these 3D representations as conditions for new video generation. However, this method is limited by the accuracy and speed of 3D reconstruction, particularly in continuously expanding large scenes where accumulated 3D reconstruction errors become intolerable. Moreover, these works focus on 3D generation and merely borrow priors from video generation models, which differs from our scope. WorldMem [42] attempts to implement memory by injecting historical frames through cross attention, and has been validated on video lengths of around $10$ seconds in Minecraft scenarios.
2.2 Context Learning for Video Generation
@@
Recently, some works [13, 12, 54] have begun to explore the role of long-context in video generation. LCT [13] performs long-context tuning on pre-trained single-shot video diffusion models to achieve consistency in multi-shot video generation. FAR [12] proposes Long-Term and Short-Term context windows to condition video generation models for long video generation. FramePack [54] introduces a hierarchical method to compress context frames into a fixed number of frames as conditioning for video generation models to achieve long video generation. However, their compression method loses too much information from temporally distant frames. In this work, we further highlight the significance of context, emphasizing that all history context serves as memory for scene-consistent long video generation.
3 Method
@@ 3.1 Preliminaries
Figure 2: Model Architecture. We concatenate the context frames to be conditioned and the predicted frames along the frame dimension. This method of injecting context is simple and effective, requiring no additional modules.
@@
Figure 3: Method Demonstration. (a) We propose Context-as-Memory, where all historical context frames serve as memory conditions in the generation of predicted frames, with Memory Retrieval extracting relevant information from all context frames. (b) Our proposed Memory Retrieval method is a search algorithm based on camera trajectories. It selects relevant frames by evaluating the overlap between camera views of different frames.
@@
Full-Sequence Text-to-Video Base Model. Our work is based on a full-sequence text-to-video model, specifically, a latent video diffusion model consisting of a causal 3D VAE [19] and a Diffusion Transformer (DiT) [29]. Each DiT block sequentially consists of spatial (2D) attention, spatial-temporal (3D) attention, cross-attention, and FFN modules. Let $\mathbf{x}$ represent a sequence of video frames, the Encoder of 3D VAE compresses it temporally and spatially to obtain the latent representation $\mathbf{z}=Encoder(\mathbf{x})$. With a temporal compression factor of $r$, the original $1+nr$ frames of $\mathbf{x}=\{x^{0},x^{1},...x^{nr}\}$ are compressed into $1+n$ latents of $\mathbf{z}=\{z^{0},z^{1},...,z^{n}\}$. During training, random Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is added to the clean latent $\mathbf{z}_{0}$ to obtain noisy latent $\mathbf{z}_{t}$ at timestep $t$. The network $\epsilon_{\phi}(\cdot)$ is trained to predict the added noise, with the following loss function:
$\mathcal{L}(\phi)=\mathbb{E}[||\boldsymbol{\epsilon}_{\phi}(\mathbf{z}_{t},\mathbf{p},t)-\boldsymbol{\epsilon}||],$ (1)
@@
where $\phi$ represents the parameters and $\mathbf{p}$ is the given text prompt. Then we use the predicted noise $\boldsymbol{\epsilon}_{\phi}$ to denoise the noisy latent. During inference, a clean latent $\mathbf{z}$ is sampled from a random Gaussian noise, then the Decoder of 3D VAE decodes it into video sequence $\mathbf{x}=Decoder(\mathbf{z})$.
@@
Camera-Conditioned Video Generation. In our work, we incorporate camera control mechanisms [41, 2] into the video generation model to implement interactive video generation. By providing camera trajectories as conditioning for video generation, we can know the camera poses of each context frame in advance. Let $\mathbf{cam}$ represent the camera poses, where $f$ denotes the total number of frames. Following the mechanism proposed in ReCamMaster [2], in order to inject $\mathbf{cam}=[R,t]\in\mathbb{R}^{f\times(3\times 4)}$, we first map it to the same dimension as the model’s feature channels through a camera encoder $\mathcal{E}_{c}(\cdot)$, followed by adding them together:
$\mathbf{F}_{i}=\mathbf{F}_{o}+\mathcal{E}_{c}(\mathbf{cam}),$ (2)
@@
where $\mathbf{F}_{o}$ is the output of spatial attention, $\mathbf{F}_{i}$ is the input of 3D attention and $\mathcal{E}_{c}(\cdot)$ is one layer of MLP with $\phi_{MLP}$ as learnable parameters. During the training of camera control, we use the original diffusion loss as follows:
$\mathcal{L}_{\mathbf{cam}}(\phi,\phi_{MLP})=\mathbb{E}[||\boldsymbol{\epsilon}_{\phi,\phi_{MLP}}(\mathbf{z}_{t},\mathbf{p},\mathbf{cam},t)-\boldsymbol{\epsilon}||].$ (3)
3.2 Context Learning Mechanism for Memory
@@
Suppose the latent of context to be conditioned is $\mathbf{z}^{c}$, and we need to learn the conditional denoiser $p(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}^{c})$. Considering that the context grows continuously during the generation process (i.e., the context is variable-length), methods designed for single-frame or fixed-length frame conditions, such as Adapter [55, 27] and channel-wise concatenation [43], are not applicable. Similar to ReCamMaster [2], we propose to inject context through concatenation along the frame dimension (shown in Fig. 2), which can flexibly support variable-length context conditions. Specifically, the clean context latents $\mathbf{z}^{c}$ participate equally with the noisy predicted latents $\mathbf{z}_{t}$ in the attention computation within DiT Blocks. During output, we only update the noisy latents $\mathbf{z}_{t}$ using the predicted noise $\boldsymbol{\epsilon}_{\phi}(\{\mathbf{z}_{t},\mathbf{z}^{c}\},\mathbf{p},t)$ while keeping the clean context latents $\mathbf{z}^{c}$ unchanged.
@@
Another challenge is how to handle positional encoding along the frame dimension in video diffusion models after context frame expansion. Since our method is based on a pre-trained full-sequence text-to-video model, to preserve the original model’s generation capability and facilitate easier adaptation to the context-conditioned generation setting, we maintain the same positional encoding for predicted latents $\mathbf{z}_{t}$ as in the pre-training phase, while assigning new positional encodings to the newly conditioned context latents $\mathbf{z}^{c}$. Our base model employs RoPE [37], which can conveniently adapt to variable-length position encodings.
3.3 Memory Retrieval
@@
Alternative method #1: random selection. A baseline randomly selects frames from context. This works well in early generation when context size is small, as adjacent frames’ natural redundancy reduces the risk of missing important information. However, with hundreds of context frames, random selection fails to identify valuable frames.
Alternative method #2: neighbor frames within a window. Another approach selects consecutive recent frames within a window near the current predicted frames. While common in existing methods [34, 5, 52], this has key limitations. First, adjacent frames’ redundancy means multiple consecutive frames add little new information beyond the most recent frame. Second, ignoring temporally distant frames prevents awareness of previously seen scenes, leading to continuous generation of new scenes and ultimately breaking scene consistency.
Alternative method #3: hierarchical compression. FramePack [54] proposes a hierarchical compression method for context frames into a minimal set (e.g., 2-3 frames). For two-frame case, it allocates space proportionally: the recent frame gets one full frame, the second recent gets half, the third gets a quarter, and so on, totaling two frames. While achieving high compression, this exponential decay significantly loses historical information. Though the authors suggest manually preserving certain key frames uncompressed, they don’t specify the selection criteria.
Our method: camera-trajectory-based search. The fundamental limitation of these methods lies in their inability to identify truly valuable frames from the large number of context frames. They either introduce many redundant frames or lose too much useful information, especially from the old frames that are temporally distant. We leverage the known camera trajectory of the context to search for valuable frames, specifically those that share high-overlap visible regions with predicted frame as shown in Fig. 3 (b).
@@
Figure 4: Examples of FOV Overlap. We simplify FOV overlap detection to checking intersections between four rays from two camera origins. A practical rule that works for most cases requires: both left and right ray pairs intersect (a, b). However, we must filter out cases where intersection points are either too near (d) or too distant (c) from cameras. While this rule may not cover all scenarios and some corner cases exist (e, f), occasional missed or incorrect candidates don’t substantially affect overall performance.

### v2

@@ Abstract.
Figure 1. Teaser Demonstration. We propose Context-as-Memory, which utilizes history context as memory to guide the generation of new frames, thereby achieving scene-consistent long video generation. This figure shows key frames in generated videos under long camera trajectories. The cameras of the key frames are marked in red. It can be observed that different frames maintain good consistency when viewing the same scene from different viewpoints.
1. Introduction
@@
Recent breakthroughs in video generation models (OpenAI, 2024; Runway, 2024; Yang et al., 2024a; Wang et al., 2025; Kong et al., 2024) have shown remarkable progress. Due to their powerful generative capabilities developed through training on large-scale real-world datasets, these models are considered to have the potential to become world models capable of modeling reality (OpenAI, 2024; Yang et al., 2024b, 2023; Qin et al., 2024). Among various research directions in this field, interactive long video generation has emerged as a crucial one since many applications, such as gaming (DeepMind, 2024a; Valevski et al., 2024; Yu et al., 2025c) and simulation (Hu et al., 2023; Gao et al., 2024; Russell et al., 2025), require interactive long video generation, where the videos are generated in a streaming manner controlled by user interactions. Recent works on long video generation (Chen et al., 2024; Song et al., 2025; Kondratyuk et al., 2023; Wang et al., 2024b; Deng et al., 2024; Gu et al., 2025; Zhang and Agrawala, 2025) have significantly facilitated research in this field.
@@
Despite these advances, current approaches still face significant challenges in terms of memory capabilities (Yu et al., 2025a, b), which refers to a model’s ability to maintain content consistency during continuous video generation, such as preserving the scene when the camera returns to a previously viewed location. Take Oasis (Decart, 2024) as an example: while it can generate lengthy Minecraft gameplay videos, even simple operations like turning left and then immediately right result in completely different scenes. This issue is prevalent across various state-of-the-art methods (Valevski et al., 2024; Kanervisto et al., 2025; Song et al., 2025; Yu et al., 2025c), suggesting that while current approaches can generate videos of extended duration, they struggle to maintain coherent long-term memory of scene content and spatial relationships.
@@
However, in our view, these methods’ limitations in memory capabilities are not surprising. This is because when generating each new video frame, these methods can only predict based on a limited number of previous frames. For instance, Diffusion Forcing (Chen et al., 2024; Song et al., 2025) can only utilize context from a fixed window of several dozen frames. While this setup works for video continuation, it fails to maintain long-term consistency. In the case of video generation, if each frame to be generated could reference all previously generated frames, the generative model could actively select and replicate relevant content from historical frames into the current frame being generated, thus it would be possible to maintain scene consistency in long videos. In other words, all previously generated context frames serve as the memory.
@@
However, the idea of ”all historical context as memory” seems intuitive but is impractical for three main reasons: (1) Including all historical frames in computation would be extremely resource-intensive. (2) Processing all historical frames is computationally wasteful since only a small fraction is relevant to the current frame generation. (3) Processing irrelevant historical frames adds noise that may hinder rather than help current frame generation. Therefore, a reasonable approach is to retrieve a small number of relevant frames from historical context as conditions for current generation, which we call ”Memory Retrieval”.
@@
  • We propose Context-as-Memory, highlighting the direct storage of frames as memory and conditioning via historical context learning for scene-consistent video generation.
  • To effectively utilize relevant history frames while minimizing costs, we design Memory Retrieval, a specialized rule-based approach using FOV overlap of camera trajectory.
@@
  • Our experiments show superior long video generation memory, significantly outperforming SOTAs and achieving effective memory even in unseen, open-domain scenarios.
2. Related Work
2.1. Interactive Long Video Generation
In following parts, we will review related work from four aspects:
Video Generation Model.
@@ 2.1. Interactive Long Video Generation
Video generation models can generate video sequences $\mathbf{x}=\{x^{0},x^{1},\dots,x^{t}\}$, where $x^{i}$ indicates the $i$-th frame. The current mainstream model architecture is based on diffusion models (Ho et al., 2020; Song et al., 2021; Song and Ermon, 2019; Lipman et al., 2022; Liu et al., 2022), which excel in generating high-quality content and have been widely adopted in video generation (OpenAI, 2024; Yang et al., 2024a; Kling, 2024; DeepMind, 2024b; Runway, 2024; Bao et al., 2024; Wang et al., 2025; Kong et al., 2024). Other alternative architectures include next-token prediction (Yan et al., 2021; Kondratyuk et al., 2023; Wang et al., 2024b) and various hybrid approaches (Chen et al., 2024; Deng et al., 2024; Li et al., 2024).
Controllable Video Generation.
@@
This task can be formulated as $p(\mathbf{x}|c)$, where $c$ represents different types of control signals. The most representative control signals include: camera motion control (Wang et al., 2024a; He et al., 2024; Fu et al., 2025; Bai et al., 2025, 2024), and agent action control in games or simulators (Valevski et al., 2024; DeepMind, 2024a; Yu et al., 2025c; Feng et al., 2024; Decart, 2024). These control signals greatly enhance user interactive experience, enabling free exploration in the created virtual worlds.
Streaming Video Generation.
@@
Streaming video generation can condition on previously generated frames to continuously generate new video frames, which can be expressed as $p(x^{0},x^{1},...,x^{n})=\prod_{i=0}^{n}p(x^{i}|x^{0},x^{1},...,x^{i-1})$, where $x^{i}$ indicates the $i$-th frame. Representative approaches include Diffusion-based methods (Chen et al., 2024; Song et al., 2025; Yu et al., 2025c; Gu et al., 2025; Zhang and Agrawala, 2025) and GPT-like next token prediction methods (Kondratyuk et al., 2023; Wang et al., 2024b; Kanervisto et al., 2025). Diffusion-based methods generally achieve higher visual quality and faster sampling speed, thus we focus on diffusion models for long video generation in this work. Although these SOTA methods generally fail to generate long videos with scene-consistent memory, instead only producing long videos with short-term continuity.
Memory Capability for Video Generation.
@@
Many related works’ demos (Valevski et al., 2024; Decart, 2024; Song et al., 2025; Kanervisto et al., 2025) have shown that current long video generation methods generally lack memory capability: while maintaining frame-to-frame continuity, the scenes continuously change. One potential approach (Yu et al., 2024a, b; Ma et al., 2024; Ren et al., 2025) is to leverage 3D reconstruction to build explicit 3D representations from generated videos, then render initial frames from these 3D representations as conditions for new video generation. However, this method is limited by the accuracy and speed of 3D reconstruction, particularly in continuously expanding large scenes where accumulated 3D reconstruction errors become intolerable. Moreover, these works focus on 3D generation and merely borrow priors from video generation models, which differs from our scope. WorldMem (Xiao et al., 2025) attempts to implement memory by injecting historical frames through cross attention, and has been validated on video lengths of around $10$ seconds in Minecraft scenarios.
2.2. Context Learning for Video Generation
@@
Recently, some works (Guo et al., 2025; Gu et al., 2025; Zhang and Agrawala, 2025) have begun to explore the role of long-context in video generation. LCT (Guo et al., 2025) performs long-context tuning on pre-trained single-shot video diffusion models to achieve consistency in multi-shot video generation. FAR (Gu et al., 2025) proposes Long-Term and Short-Term context windows to condition video generation models for long video generation. FramePack (Zhang and Agrawala, 2025) introduces a hierarchical method to compress context frames into a fixed number of frames as conditioning for video generation models to achieve long video generation. However, their compression method loses too much information from temporally distant frames. In this work, we further highlight the significance of context, emphasizing that all history context serves as memory for scene-consistent long video generation.
3. Method
@@ 3.1. Preliminaries
Figure 2. Model Architecture. We concatenate the context to be conditioned and the predicted frames along the frame dimension. This method of injecting context is simple and effective, requiring no additional modules.
@@
Figure 3. Method Demonstration. (a) We propose Context-as-Memory, where all historical context frames serve as memory conditions in the generation of predicted frames, with Memory Retrieval extracting relevant information from all context frames. (b) Our proposed Memory Retrieval method is a search algorithm based on camera trajectories. It selects relevant frames by evaluating the overlap between camera views of different frames.
Full-Sequence Text-to-Video Base Model.
@@
Our work is based on a full-sequence text-to-video model, specifically, a latent video diffusion model consisting of a causal 3D VAE (Kingma et al., 2013) and a Diffusion Transformer (DiT) (Peebles and Xie, 2023). Each DiT block sequentially consists of spatial (2D) attention, spatial-temporal (3D) attention, cross-attention, and FFN modules. Let $\mathbf{x}$ represent a sequence of video frames, the Encoder of 3D VAE compresses it temporally and spatially to obtain the latent representation $\mathbf{z}=Encoder(\mathbf{x})$. With a temporal compression factor of $r$, the original $1+nr$ frames of $\mathbf{x}=\{x^{0},x^{1},...x^{nr}\}$ are compressed into $1+n$ latents of $\mathbf{z}=\{z^{0},z^{1},...,z^{n}\}$. During training, random Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is added to the clean latent $\mathbf{z}_{0}$ to obtain noisy latent $\mathbf{z}_{t}$ at timestep $t$. The network $\epsilon_{\phi}(\cdot)$ is trained to predict the added noise, with the following loss function:
[…]
@@
where $\phi$ represents the parameters and $\mathbf{p}$ is the given text prompt. Then we can use the predicted noise $\boldsymbol{\epsilon}_{\phi}$ to denoise the noisy latent. During inference, a clean latent $\mathbf{z}$ can be sampled from a randomly sampled Gaussian noise, then the Decoder of 3D VAE decodes it into video sequence $\mathbf{x}=Decoder(\mathbf{z})$.
Camera-Conditioned Video Generation.
@@
In our work, we incorporate camera control mechanisms (Wang et al., 2024a; Bai et al., 2025) into the video generation model to implement interactive video generation. By providing camera trajectories as conditioning for video generation, we can know the camera poses of each context frame in advance. Let $\mathbf{cam}$ represent the camera poses, where $f$ denotes the total number of frames. Following the mechanism proposed in ReCamMaster (Bai et al., 2025), in order to inject $\mathbf{cam}=[R,t]\in\mathbb{R}^{f\times(3\times 4)}$, we first map it to the same dimension as the model’s feature channels through a camera encoder $\mathcal{E}_{c}(\cdot)$, followed by adding them together:
(2) $\mathbf{F}_{i}=\mathbf{F}_{o}+\mathcal{E}_{c}(\mathbf{cam}),$
@@
where $\mathbf{F}_{o}$ is the output of spatial attention module, $\mathbf{F}_{i}$ is the input of 3D attention module and $\mathcal{E}_{c}(\cdot)$ is one layer of MLP with $\phi_{MLP}$ as learnable parameters. During the training of camera control, we use the original diffusion loss as follows:
[…]
3.2. Context Frames Learning Mechanism for Memory
@@
Suppose the latent of context that needs to be conditioned is $\mathbf{z}^{c}$, and we need to learn the conditional denoiser $p(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}^{c})$. Considering that the context grows continuously during the generation process (i.e., the context is variable-length), methods designed for single-frame or fixed-length frame conditions, such as Adapter (Zhang et al., 2023; Mou et al., 2024) and channel-wise concatenation (Xing et al., 2023), are not applicable. Similar to ReCamMaster (Bai et al., 2025), we propose to inject context through concatenation along the frame dimension (shown in Fig. 2), which can flexibly support variable-length context conditions. Specifically, the clean context latents $\mathbf{z}^{c}$ participate equally with the noisy predicted latents $\mathbf{z}_{t}$ in the attention computation within DiT Blocks. During output, we only update the noisy latents $\mathbf{z}_{t}$ using the predicted noise $\boldsymbol{\epsilon}_{\phi}(\{\mathbf{z}_{t},\mathbf{z}^{c}\},\mathbf{p},t)$ while keeping the clean context latents $\mathbf{z}^{c}$ unchanged.
@@
Another challenge is how to handle positional encoding along the frame dimension in video diffusion models after context frame expansion. Since our method is based on a pre-trained full-sequence text-to-video model, to preserve the original model’s generation capability and facilitate easier adaptation to the context-conditioned generation setting, we maintain the same positional encoding for predicted latents $\mathbf{z}_{t}$ as in the pre-training phase, while assigning new positional encodings to the newly conditioned context latents $\mathbf{z}^{c}$. Our base model employs RoPE (Su et al., 2024), which can conveniently adapt to variable-length position encodings.
3.3. Memory Retrieval
@@
Alternative method #1: random selection.
A baseline randomly selects frames from context. This works well in early generation when context size is small, as adjacent frames’ natural redundancy reduces the risk of missing important information. However, with hundreds of context frames, random selection fails to identify valuable frames.
Alternative method #2: neighbor frames within a window.
Another approach selects consecutive recent frames within a window near the current predicted frames. While common in existing methods (Song et al., 2025; Decart, 2024; Yu et al., 2025c), this has key limitations. First, adjacent frames’ redundancy means multiple consecutive frames add little new information beyond the most recent frame. Second, ignoring temporally distant frames prevents awareness of previously seen scenes, leading to continuous generation of new scenes and ultimately breaking scene consistency.
Alternative method #3: hierarchical compression.
FramePack (Zhang and Agrawala, 2025) proposes a hierarchical compression method for context frames into a minimal set (e.g., 2-3 frames). For two-frame compression, it allocates space proportionally: the most recent frame gets one full frame, the second most recent gets half, the third gets a quarter, and so on, totaling two frames. While achieving high compression, this exponential decay significantly loses historical information. Though the authors suggest manually preserving certain key frames uncompressed, they don’t specify the selection criteria.
Our method: camera-trajectory-based search.
The fundamental limitation of these methods lies in their inability to identify truly valuable frames from the large number of context frames. They either introduce many redundant frames or lose too much useful information, especially from the old frames that are temporally distant. We leverage the known camera trajectory of the context to search for valuable frames, specifically those that share high-overlap visible regions with predicted frame as shown in Fig. 3 (b).
@@
Figure 4. Examples of FOV Overlap. We simplify FOV overlap detection to checking intersections between four rays from two camera origins. A practical rule that works for most cases requires: both left and right ray pairs intersect (a, b). However, we must filter out cases where intersection points are either too near (d) or too distant (c) from cameras. While this rule may not cover all scenarios and some corner cases exist (e, f), occasional missed or incorrect candidates don’t substantially affect overall performance.
The second question is how to determine co-visibility between two frames given their camera poses. We attempt to determine this by checking if there is an overlapping region between the fan-shaped areas corresponding to the Fields of View (FOV) of the two cameras. Specifically, since we restrict camera movement to the XY plane, we only need to consider the left and right rays shooting from each camera’s origin. By checking the intersection of these four rays from two cameras, we can quickly determine the FOV overlap as shown in Fig. 4. Additionally, we calculate the distance between the predicted frame’s camera and the calculated intersection points to eliminate cases where the cameras are too far apart (which typically indicates no actual overlap or very small overlap). This FOV overlap detection is not perfect, as it may fail in cases with occlusions. However, this method effectively reduces the number of candidate context frames.
Implementation details in training and inference.
Assume the maximum number of retrieved context frames is $k$. During training, we read a long ground truth video (containing thousands of frames) and randomly select a segment as the sequence to be predicted. We then apply our Memory Retrieval method to select k1k-1 context frames from the remaining frames. The overlapping relationships between frames have been pre-computed, eliminating the need for repeated calculations. The first frame of the prediction sequence is also included as an additional context frame to ensure video continuity. Additionally, there is a $10\%$ probability during training that only the recent context frame is used, simulating the beginning of long video generation where no context frames are available. During inference, for each video segment to be predicted, we search k1k-1 context frames from the previously generated frames using FOV-based Memory Retrieval and add the most recently generated frame to the context. The training and inference procedures are outlined in Algorithm 1 and 2, respectively.